# Inverse Scaling: When Bigger Isn’t Better

### v1

@@ 1 Introduction
However, the task of predicting human-written text is importantly different from many real-world tasks, and we hypothesize that text prediction actively trains LMs to behave in undesirable ways for many tasks. This paper focuses on inverse scaling (Lin et al., 2022): a phenomenon where task performance gets worse as loss on the original training objective gets better. When used to perform tasks that they were not explicitly trained on, like question-answering or sentiment analysis, LMs perform well only insofar as the training objective encourages theimodel to generalize well to these tasks. This dynamic leaves open the possibility that bad performance on some tasks is actively incentivized by the objective. Given the widespread adoption of LM training in state-of-the-art systems, it is critical to identify cases of inverse scaling tasks in order to refine our understanding of what LM training teaches models and where it fails. A better understanding of LM training in turn can help us develop mitigation strategies for the issues found, e.g., by fixing those failures in later stages of training (Ouyang et al., 2022) or by improving the pretraining process (Korbak et al., 2023).
To this end, we ran a public contest to collect examples of inverse scaling (§2). We evaluated submissions in zero-shot (no examples provided in the input) and few-shot (a few examples provided) settings across model series from OpenAI, Anthropic, and DeepMind, covering over 5 orders of magnitude: $10^{18}$ to $10^{23}$ training FLOPs. [^1] Training FLOPs measure the amount of compute used during LM pretraining and correlate with model size, training time, and data quantity. We focus on training FLOPs rather than model parameters because training compute is a better proxy for LM performance (Hoffmann et al., 2022). We also show results on models with and without instruction-tuning (Ouyang et al., 2022; Bai et al., 2022), to understand the extent to which training models to follow instructions helps to mitigate undesirable behaviors in LMs. The contest attracted 99 submissions over two rounds, and we awarded prizes to 11 submissions that appeared to robustly demonstrate inverse scaling on the models we evaluated, including several held-out model series. Many of the instances of inverse scaling we found are straightforward tasks that humans perform with ease (verified with crowdworker annotation).
@@ 2.2 Submission Format and Metrics
  • Classification Loss (classification). This metric can be used for standard classification tasks, for example when testing how well a model can choose the correct response. Each class could consist of multiple tokens, so we used the probability of the full token sequences (renormalized to sum to 1) to compute the classification loss, by evaluating the average negative log-probability of the correct response.
@@ 3 Inverse Scaling Prize Tasks
In Table 1, we provide an overview of all winning tasks, including the total number of examples provided, as well as human agreement with the task labels on a random sample of at least 50 examples. [^6] Human validation was done by Surge AI. For tasks that were submitted in multiple parts, we took 50 examples from each part and averaged the agreement scores. For tasks with 10 or more parts (like Sig Figs), we manually grouped similar parts together and took 50 samples from each group. In Round 1, we received 50 submissions and awarded 4 third prizes. In Round 2, we received 49 submissions and awarded 7 additional third prizes, as well as accepting updates to the datasets of two Round 1 winners. We awarded 11 third prizes in total (more than the initially planned 10). We did not award any grand or second prizes because no submitted tasks met our criteria for those prizes (see §3.5 for more discussion). We release the data at https://inversescaling.com/data under a CC BY 4.0 license. [^7] https://creativecommons.org/licenses/by/4.0/
@@ 3.1.1 Resisting Correction
Figure 2 (left) shows the scaling behavior of this task. Small LMs will typically correctly repeat the word, while larger LMs fail at this task more often. One hypothesis for this behavior is that larger LMs have a stronger prior on grammatical sentences, and so will have a harder time overriding this prior, even when explicitly directed to do so. There are signs of U-shaped scaling on PaLM, OPT, and the DeepMind models (Gopher and Chinchilla). However, only Chinchilla has better accuracy on its smallest model than on its largest model.
@@ 3.1.4 Prompt Injection
Figure 4 shows the scaling behavior of this task. At small scales, models have high loss, which drops by around $10^{21}$ FLOPs. After this point, loss gets worse for all model series. Importantly, the scaling trends we observe on this task are in the opposite direction to the U-shaped trends discussed in Wei et al. (2022a), with trends reversing for the worse at large scales, which we call inverted-U scaling. Thus, many model series have an inverted-U scaling trend, and others that start at higher FLOP counts show inverse scaling.
@@ 3.4.1 Hindsight Neglect
This task is important as it demonstrates that correctly-labeled few-shot examples can still cause the model to answer incorrectly by demonstrating a spurious correlation (in this case whether the outcome matched the expected value). Few-shot learning is a common and natural way to specify tasks for LMs to perform, and it is infeasible to demonstrate intended behavior in all situations with the chosen examples. Underspecification in the task could in turn lead to goal misgeneralization (Shah et al., 2022), where the LM competently performs a task that is compatible with the given few-shot examples but was not intended.
@@
Figure 7 (left) shows the scaling behavior of this task. All models start out around random performance, falling off to below random performance at around $10^{22}$ training FLOPs. GPT-4 performs well on this task, [^11] GPT-4 performance is taken from OpenAI (2023), but it is unclear whether the model used there has been trained with RLHF or not. PaLM shows strong U-shaped scaling, and there are some signs of U-shaped scaling trends on OPT and GPT-3 FeedME, but inverse scaling is strong on the DeepMind and Anthropic models.
@@ 4.1 Language Model Evaluation Suites
Several multi-task benchmarks have been created that attempt to provide an overall picture of the ability of LMs. GLUE (Wang et al., 2018) and its successor SuperGLUE (Wang et al., 2019) are benchmarks of diverse tasks aimed at testing natural language understanding. Human performance has been met (or exceeded) on both GLUE and SuperGLUE. MMLU (Hendrycks et al., 2021) is a benchmark of 57 tasks covering different topics designed to test the breadth and depth of the world knowledge and problem-solving ability of an LM. MMLU focuses on tasks that are expected to improve with scale and thus does not include coverage or discussion of inverse scaling. BIG-Bench (Srivastava et al., 2022) is a large collection of more than 200 tasks sourced from the LM research community. As discussed in §4.2.3 below, BIG-Bench contains some tasks that demonstrate inverse scaling. Some BIG-Bench tasks also show U-shaped scaling, as discussed in Wei et al. (2022a). HELM (Liang et al., 2022) is a living benchmark intended to holistically evaluate the capabilities and limitations LMs. They evaluate 30 LMs across 42 use cases of LMs. HELM discusses trends with model scale but does not mention the possibility of inverse scaling or have evaluations focused on inverse scaling.
@@ D.1 Inverse Scaling Prize Rubric
Inverse Scaling Strength
How straight and steep is the inverse scaling trend on public models?
Shows flat, very bumpy, or standard scaling.
Shows approximately monotonic inverse scaling.
Shows a clear, strictly monotonic inverse scaling trend.
Inverse Scaling Generality
Do different models all show inverse scaling?
No inverse scaling on private models.
Shows inverse scaling on some public and some private models.
Shows inverse scaling across all public and private models tested.
Task Importance
Is the task important to the safe and responsible use of LMs, or for shedding light on where LMs fail? How strong are the arguments?
Weak. No users or third parties would be harmed, and the task does not shed light on where LMs fail.
Fairly convincing. Some LM users or third parties would be harmed by the discovered behavior, or the task sheds light on where LMs fail.
Very convincing. Significant implications for how LM research or deployment will need to be developed to be reliably safe and effective.
Novelty and Surprisingness
Is inverse scaling on the task novel (not shown in prior work) and surprising?
Not novel or surprising.
Novel and somewhat surprising.
Novel and surprising, teaching us something new about LMs.
Task Coverage
Are the examples fully representative of the described task?
Examples only cover a special subcategory or phrasing of the described task. There’s no evidence of inverse scaling on other subcategories or phrasings.
Examples cover different subcategories and phrasings for the described task.
Examples cover almost all important task subcategories and phrasings, suggesting robust inverse scaling on the described task.
Reproducibility
Does inverse scaling appear to occur if we reproduce the task based on its description?
No. The particular examples submitted may have been over-optimized for inverse scaling, to the extent that the examples are unrepresentative of the described task.
Yes, but to a lesser extent.
Yes, to a similar or stronger extent.

### v2

@@ 1 Introduction
However, the task of predicting human-written text is importantly different from many real-world tasks, and we hypothesize that text prediction actively trains LMs to behave in undesirable ways for many tasks. This paper focuses on inverse scaling (Lin et al., 2022): a phenomenon where task performance gets worse as loss on the original training objective gets better. When used to perform tasks that they were not explicitly trained on, like question-answering or sentiment analysis, LMs perform well only insofar as the training objective encourages the model to generalize well to these tasks. This dynamic leaves open the possibility that bad performance on some tasks is actively incentivized by the objective. Given the widespread adoption of LM training in state-of-the-art systems, it is critical to identify cases of inverse scaling tasks in order to refine our understanding of what LM training teaches models and where it fails. A better understanding of LM training in turn can help us develop mitigation strategies for the issues found, e.g., by fixing those failures in later stages of training (Ouyang et al., 2022) or by improving the pretraining process (Korbak et al., 2023).
To this end, we ran a public contest to collect examples of inverse scaling (§2). We evaluated submissions in zero-shot (no examples provided in the input) and few-shot (a few examples provided) settings across model series from OpenAI, Anthropic, and DeepMind, covering over 5 orders of magnitude: $10^{18}$ to $10^{23}$ training FLOPs. [^1] Training FLOPs measure the amount of compute used during LM pretraining and correlate with model size, training time, and data quantity. We focus on training FLOPs rather than model parameters because training compute is a better proxy for LM performance (Hoffmann et al., 2022). However, since different model families have different ratios of model size to data quantity, comparisons based on FLOPs between model families can be misleading. We also show results on models with and without instruction-tuning (Ouyang et al., 2022; Bai et al., 2022), to understand the extent to which training models to follow instructions helps to mitigate undesirable behaviors in LMs. The contest attracted 99 submissions over two rounds, and we awarded prizes to 11 submissions that appeared to robustly demonstrate inverse scaling on the models we evaluated, including several held-out model series. Many of the instances of inverse scaling we found are straightforward tasks that humans perform with ease (verified with crowdworker annotation).
@@
To summarize our contributions:
  • We discovered 11 cases of inverse scaling in LMs through a public contest (§2).
  • We identified possible causes of inverse scaling behavior, and through a systematic examination, classified the 11 cases of inverse scaling into categories that offer an explanation about the cause (§3).
  • We collected instances of inverse scaling in the literature and show how they fit into our explanation of why inverse scaling happens in LMs (§4.2).
@@ 2.2 Submission Format and Metrics
  • Classification Loss (classification). This metric can be used for standard classification tasks, for example when testing how well a model can choose the correct response. Each class could consist of multiple tokens, so we used the probability of the full token sequences (renormalized to sum to 1) to compute the classification loss, by evaluating the average negative log-probability of the correct response. Normalization is performed over the options given, rather than over all sequences.
@@ 3 Inverse Scaling Prize Tasks
In Table 1, we provide an overview of all winning tasks, including the total number of examples provided, as well as human agreement with the task labels on a random sample of at least 50 examples. The purpose of the human agreement scores is to ensure that the labels given by task authors represent the answer that humans think is correct. This is because it is easy to construct an inverse scaling task if the answers were allowed to be incorrect: a dataset of simple True/False questions with the True and False labels switched would almost certainly show inverse scaling. Contractors were allowed to use the internet to help them come to a final answer to each question, to reduce mistakes or a lack of knowledge affecting the scores. [^6] Human validation was done by Surge AI. For tasks that were submitted in multiple parts, we took 50 examples from each part and averaged the agreement scores. For tasks with 10 or more parts (like Sig Figs), we manually grouped similar parts together and took 50 samples from each group.
In Round 1, we received 50 submissions and awarded 4 third prizes. In Round 2, we received 49 submissions and awarded 7 additional third prizes, as well as accepting updates to the datasets of two Round 1 winners. We awarded 11 third prizes in total (more than the initially planned 10). We did not award any grand or second prizes because no submitted tasks met our criteria for those prizes (see §3.5 for more discussion). We release the data at https://inversescaling.com/data under a CC BY 4.0 license. [^7] https://creativecommons.org/licenses/by/4.0/
@@ 3.1.1 Resisting Correction
Figure 2 (left) shows the scaling behavior of this task. Small LMs will typically correctly repeat the word, while larger LMs fail at this task more often. One hypothesis for this behavior is that larger LMs have a stronger prior on grammatical sentences, and so will have a harder time overriding this prior, even when explicitly directed to do so. There are signs of U-shaped scaling on PaLM, OPT, and the DeepMind models (Gopher and Chinchilla). However, only Chinchilla has better accuracy on its largest model than on its smallest model.
@@ 3.1.2 Memo Trap
One limitation of this task is that LMs could be intending a multi-word completion which starts with the memorized word, but follows it with a novel ending. An improved version of this task might add a period to the end of the options to indicate that the sentence ends after that word.
@@ 3.1.4 Prompt Injection
Figure 4 shows the scaling behavior of this task. [^10] Figure 4 uses negated loss on the y-axis. See Figure 10 in Appendix E for scaling behavior when probability is used on the y-axis instead. At small scales, models have high loss, which drops by around $10^{21}$ FLOPs. After this point, loss gets worse for all model series. Importantly, the scaling trends we observe on this task are in the opposite direction to the U-shaped trends discussed in Wei et al. (2022a), with trends reversing for the worse at large scales, which we call inverted-U scaling. Thus, many model series have an inverted-U scaling trend, and others that start at higher FLOP counts show inverse scaling.
@@
Prompt Injection uses a metric only based on loss, unlike the other tasks, which use accuracy scores. This means that for this task, one mechanism through which inverse scaling could occur is increasing model confidence on incorrect answers.
@@ 3.2.1 Modus Tollens
After preparation of the paper, some grammatical errors were discovered that affected less than 10% of examples. Removing the affected examples had very little effect on the results and so the analysis is unchanged, but we include a version of the plot with those examples removed in Appendix E.
@@ 3.4.1 Hindsight Neglect
This task is important as it demonstrates that correctly-labeled few-shot examples can still cause the model to answer incorrectly by demonstrating a spurious correlation (in this case whether the outcome matched the expected value). Few-shot learning is a common and natural way to specify tasks for LMs to perform, and it is infeasible to demonstrate intended behavior in all situations with the chosen examples. Underspecification in the task could in turn lead to goal misgeneralization (Shah et al., 2022), where the LM competently performs a task that is compatible with the given few-shot examples but is not the intended task.
@@
Figure 7 (left) shows the scaling behavior of this task. All models start out around random performance, falling off to below random performance at around $10^{22}$ training FLOPs. GPT-4 performs well on this task, [^12] GPT-4 performance is taken from OpenAI (2023), but it is unclear whether the model used there has been trained with RLHF or not. PaLM shows strong U-shaped scaling, and there are some signs of U-shaped scaling trends on OPT and GPT-3 FeedME, but inverse scaling is strong on DeepMind and Anthropic models.
@@ 4.1 Language Model Evaluation Suites
Several multi-task benchmarks have been created that attempt to provide an overall picture of the ability of LMs. GLUE (Wang et al., 2018) and its successor SuperGLUE (Wang et al., 2019) are benchmarks of diverse tasks aimed at testing natural language understanding. Human performance has been met (or exceeded) on both GLUE and SuperGLUE. MMLU (Hendrycks et al., 2021) is a benchmark of 57 tasks covering different topics designed to test the breadth and depth of the world knowledge and problem-solving ability of an LM. MMLU focuses on tasks that are expected to improve with scale and thus does not include coverage or discussion of inverse scaling. BIG-Bench (Srivastava et al., 2022) is a large collection of more than 200 tasks sourced from the LM research community. As discussed in §4.2.3 below, BIG-Bench contains some tasks that demonstrate inverse scaling, but does not actively solicit inverse scaling tasks, and does not discuss potential causes. Some BIG-Bench tasks also show U-shaped scaling, as discussed in Wei et al. (2022a). HELM (Liang et al., 2022) is a living benchmark intended to holistically evaluate the capabilities and limitations LMs. They evaluate 30 LMs across 42 use cases of LMs. HELM discusses trends with model scale but does not mention the possibility of inverse scaling or have evaluations focused on inverse scaling.
@@ D.1 Inverse Scaling Prize Rubric
Inverse Scaling Strength How straight and steep is the inverse scaling trend on public models? Shows flat, very bumpy, or standard scaling. Shows approximately monotonic inverse scaling. Shows a clear, strictly monotonic inverse scaling trend.
Inverse Scaling Generality Do different models all show inverse scaling? No inverse scaling on private models. Shows inverse scaling on some public and some private models. Shows inverse scaling across all public and private models tested.
Task Importance Is the task important to the safe and responsible use of LMs, or for shedding light on where LMs fail? How strong are the arguments? Weak. No users or third parties would be harmed, and the task does not shed light on where LMs fail. Fairly convincing. Some LM users or third parties would be harmed by the discovered behavior, or the task sheds light on where LMs fail. Very convincing. Significant implications for how LM research or deployment will need to be developed to be reliably safe and effective.
Novelty and Surprisingness Is inverse scaling on the task novel (not shown in prior work) and surprising? Not novel or surprising. Novel and somewhat surprising. Novel and surprising, teaching us something new about LMs.
Task Coverage Are the examples fully representative of the described task? Examples only cover a special subcategory or phrasing of the described task. There’s no evidence of inverse scaling on other subcategories or phrasings. Examples cover different subcategories and phrasings for the described task. Examples cover almost all important task subcategories and phrasings, suggesting robust inverse scaling on the described task.
Reproducibility Does inverse scaling appear to occur if we reproduce the task based on its description? No. The particular examples submitted may have been over-optimized for inverse scaling, to the extent that the examples are unrepresentative of the described task. Yes, but to a lesser extent. Yes, to a similar or stronger extent.
@@ D.5 Submission Assessment
Appendix E Additional Plots
E.1 Modus Tollens Corrected Plot
After preparation of the paper, some grammatical errors were discovered that affected less than 10% of examples from the Modus Tollens task. Removing the affected examples had very little effect on the results and so the analysis is unchanged, but we include a fixed version of the plot in Figure 9.
Figure 9: Scaling behavior for a version of Modus Tollens (§3.2.1) with examples containing grammatical errors removed.
E.2 Prompt Injection Probability Plot
Prompt Injection (§3.1.4) uses the (sequence prob) metric, which does not have an accuracy score. It can be difficult to interpret the results when only looking at loss, so we include a plot of average probability in Figure 10.
Figure 10: Scaling behavior for probability on Prompt Injection (§3.1.4).